OpenStudy (anonymous):

If the factors of a polynomial are x+2 and x+6, what values of x make that polynomial 0? A) -2 and -6 B) 1 and 2 C) 2 and 6 D) Cannot be determined

OpenStudy (anonymous):

umm A

jimthompson5910 (jim_thompson5910):

If the factors are x+2 and x+6, then we can say (x+2)(x+6) = 0

jimthompson5910 (jim_thompson5910):

Which breaks up into x+2 = 0 or x+6 = 0

OpenStudy (anonymous):

its A!!

OpenStudy (anonymous):

Oh okay. Thanks Katie. :)

OpenStudy (anonymous):

aint this the second time ive saved u lol or was that some one else